### v1

# AHCPTQ: Accurate and Hardware-Compatible Post-Training Quantization for Segment Anything Model

Abstract
The Segment Anything Model (SAM) has demonstrated strong versatility across various visual tasks. However, its large storage requirements and high computational cost pose challenges for practical deployment. Post-training quantization (PTQ) has emerged as an effective strategy for efficient deployment, but we identify two key challenges in SAM that hinder the effectiveness of existing PTQ methods: the heavy-tailed and skewed distribution of post-GELU activations, and significant inter-channel variation in linear projection activations. To address these challenges, we propose AHCPTQ, an accurate and hardware-efficient PTQ method for SAM. AHCPTQ introduces hardware-compatible Hybrid Log-Uniform Quantization (HLUQ) to manage post-GELU activations, employing log2 quantization for dense small values and uniform quantization for sparse large values to enhance quantization resolution. Additionally, AHCPTQ incorporates Channel-Aware Grouping (CAG) to mitigate inter-channel variation by progressively clustering activation channels with similar distributions, enabling them to share quantization parameters and improving hardware efficiency. The combination of HLUQ and CAG not only enhances quantization effectiveness but also ensures compatibility with efficient hardware execution. For instance, under the W4A4 configuration on the SAM-L model, AHCPTQ achieves 36.6% mAP on instance segmentation with the DINO detector, while achieving a 7.89×7.89$\times$7.89 × speedup and 8.64×8.64$\times$8.64 × energy efficiency over its floating-point counterpart in FPGA implementation. Code will be available upon acceptance.
1 Introduction
The Segment Anything Model (SAM) [13] is a powerful tool for promptable segmentation, demonstrating strong zero-shot performance across diverse visual domains and broad applicability in real-world scenarios [2, 41, 37, 29, 32]. However, its large-scale parameters, substantial storage demands, and high computational costs pose significant challenges for deployment on edge devices [11, 42, 18, 27].
(a) Challenge 1: Heavy-tailed and skewed post-GELU activations (from the 10th block of the image encoder).
(b) Challenge 2: Inter-channel variation for channel indices 50 to 100 (from the first block of the mask decoder).
Figure 1: Challenges in SAM quantization (Data obtained from the SAM-B model with the YOLOX detector).
To tackle this challenge, model quantization has been widely adopted to replace floating-point weights and activations with low-bit representations, reducing storage overhead and enabling efficient integer-based computations, making it well-suited for edge devices with constrained resources [14]. One prominent approach is Quantization-Aware Training (QAT), which incorporates quantization effects during training, allowing the model to adapt to quantized weights and activations [5, 17, 15, 8, 24]. However, applying QAT to SAM is impractical due to the computational expense of SAM training, which relies on a co-developed data engine and consequently utilizes the SA-1B dataset comprising 1.1B masks and 11M images [13]. As a more practical alternative, Post-Training Quantization (PTQ) has gained increasing attention. PTQ requires only a small calibration dataset, significantly reducing data and computational demands while maintaining competitive accuracy [19, 22, 26, 6, 33, 45, 46, 44, 4].
PTQ4SAM [28] has recently demonstrated the feasibility of applying PTQ to SAM [25, 34], utilizing equivalent sign transformations and adaptive resolution quantization to accommodate its unique activation distributions. Despite these advancements, our study reveals that existing PTQ methods suffer from quantization collapse at 4-bit and severe accuracy degradation at 5-bit, limiting their practical utility for ultra-low-bit deployment. As presented in Table 1, PTQ4SAM achieves only 3.8% mAP at 4-bit and 18.4% mAP at 5-bit quantization on the YOLOX detector with the SAM-B model, compared to 37.2% and 40.4% mAP in floating-point, respectively. Such substantial accuracy loss significantly hampers the feasibility of deploying SAM on edge devices, motivating the need for more effective ultra-low-bit quantization techniques.
In this paper, we identify two additional challenges that limit PTQ performance on the SAM model. First, as illustrated in Fig. 1(a), post-GELU activations exhibit an extremely imbalanced distribution, where numerous small values are densely clustered within a narrow range, while sparse large values extend across a wide range. This imbalance poses a significant challenge for traditional hardware-friendly uniform and log2 quantizers. Although non-uniform quantizers [35, 3, 9] can better adapt to this distribution, they introduce substantial challenges in hardware implementation. Second, activations in Query/Key/Value linear projections and Linear-1 of MLP exhibit high inter-channel variance, as exemplified in Fig. 1(b). Such variations render per-tensor quantization ineffective—a small scale factor results in large quantization errors for channels with wide dynamic ranges, while a large scale factor zeros out activations in channels with small dynamic ranges. Although per-channel quantization improves accuracy, it introduces deployment inefficiencies due to the high overhead of storing and transferring quantization parameters. Addressing these challenges requires a quantization strategy that both closely aligns with activation distributions and remains compatible with efficient hardware deployment.
To overcome the challenges in PTQ for SAM, we present AHCPTQ, a novel approach that enhances quantization effectiveness while ensuring hardware compatibility. AHCPTQ addresses the two identified challenges through Hybrid Log-Uniform Quantization (HLUQ) and Channel-Aware Grouping (CAG), respectively. HLUQ innovatively employs log2 quantization for densely clustered small values and uniform quantization for sparse but widely distributed large values, effectively adapting to the heavy-tailed and skewed distribution of post-GELU activations while maintaining hardware-efficient implementation. Meanwhile, CAG selectively clusters channels with similar statistical properties, assigning a shared quantization parameter to each group. This method achieves accuracy comparable to per-channel quantization while reducing on-chip register overhead by 99.7% and enabling efficient storage of quantization parameters within a compact set of on-chip registers. As a result, AHCPTQ significantly reduces accuracy loss at 5-bit PTQ and, to the best of our knowledge, is the first to enable 4-bit PTQ on SAM. Our main contributions are summarized as follows:
  • We identify two key challenges limiting PTQ performance on SAM and propose AHCPTQ, a framework that integrates CAG and HLUQ to achieve a balance between quantization effectiveness and hardware feasibility.
  • Extensive experiments show that AHCPTQ consistently surpasses previous PTQ methods, setting a new state-of-the-art in SAM quantization. Under the W4A4 configuration, AHCPTQ achieves 36.6% mAP on instance segmentation with the DINO detector on SAM-L, representing a substantial improvement over PTQ4SAM, which fails to function under the same conditions.
  • We further develop an FPGA-based accelerator to evaluate AHCPTQ’s hardware efficiency. Our results indicate that AHCPTQ delivers a 7.89×7.89$\times$7.89 × speedup and 8.64×8.64$\times$8.64 × energy efficiency improvement over floating-point implementations, demonstrating superior resource utilization.
2 Related Works
PTQ utilizes a small calibration dataset to determine quantization parameters, enabling rapid deployment on edge devices without requiring extensive retraining. This paper focuses on learning-based PTQ methods, which leverage optimization techniques to derive optimal quantization parameters, often outperforming static PTQ methods. AdaRound [30] identifies the sensitivity of weight rounding and introduces an optimization technique to reduce overall model loss. BRECQ [16] employs block reconstruction to strike a balance between cross-layer dependency and generalization error. QDrop [38] integrates dropout into the reconstruction process to improve the flatness of the optimized models. Despite their success, these methods are primarily designed for CNN-based models and face challenges when applied to Transformer architectures. In the domain of Transformer-based models, FQ-ViT [22] improves granularity using powers-of-two scale and Log-Int-Softmax while maintaining hardware efficiency. RepQ-ViT [20] eliminates parameter overhead in per-channel quantization by conducting reparameterization techniques to post-LayerNorm activations. ERQ [46, 45] minimizes quantization errors via ridge regression, while PTQ4ViT [39] employs a twin uniform quantizer to efficiently handle post-Softmax and post-GELU activations. While these approaches offer insights for mitigating challenges in SAM, they are not easily adaptable to learning-based PTQ, and the mask decoder architecture of SAM restricts reparameterization techniques. PTQ4SAM [28], the first PTQ method specialized for SAM, introduces bimodal integration and adaptive log quantization to address unique distribution challenges while maintaining hardware efficiency. However, it fails to achieve 4-bit quantization functionality, motivating us to explore additional constraints and propose novel strategies to further improve performance.
3 Methodology
3.1 Preliminaries
3.1.1 Quantizer
Quantization discretizes continuous values into low-bit representations, enabling efficient computation and reducing memory usage. Two widely adopted quantization schemes are uniform quantization and log2 quantization. The uniform quantizer divides the input range into equally spaced intervals, mapping each value to the closest quantized level:
$x_{q}=\text{clamp}\left(\left\lfloor\frac{x}{s}\right\rceil+z,0,2^{k}-1\right).$ (1)
$x\approx\hat{x}=s\cdot(x_{q}-z).$ (2)
[…]
$x_{q}=\text{clamp}\left(\left\lfloor-\log_{2}\frac{x}{s}\right\rceil,0,2^{k}-1\right)$ (3)
$x\approx\hat{x}=s\cdot 2^{-x_{q}}$ (4)
Log2 quantization is particularly beneficial for hardware, as it enables multiplications to be replaced by bit shifts, improving computational speed and energy efficiency.
3.1.2 Quantization Granularity
Quantization operates at varying levels of granularity, introducing a trade-off between computational efficiency and quantization effectiveness. The two most common approaches are per-tensor and per-channel quantization. Per-Tensor quantization employs a single scale and zero-point across an entire weight or activation tensor, reducing computational complexity and memory overhead. However, it struggles with large inter-channel variations, leading to suboptimal quantization performance. Per-Channel quantization assigns individual quantization parameters to each output channel, effectively mitigating distribution variance across channels. However, it requires storing more quantization parameters, increasing memory usage and data transfer costs.
3.1.3 Block-Wise Reconstruction
We employ block-wise reconstruction [38], as adopted in PTQ4SAM [28], to mitigate the quantization-induced error in weight and activation quantization by minimizing the mean squared error:
$\mathcal{L}=\|\mathbf{O}_{\mathcal{B}}-\hat{\mathbf{O}}_{\mathcal{B}}\|_{2}^{2},$ (5)
where $\mathbf{O}_{\mathcal{B}}$ and $\hat{\mathbf{O}}_{\mathcal{B}}$ represent the floating-point and quantized outputs of the $\mathcal{B}$-th block, respectively.
3.2 Challenges of SAM Quantization
Recent PTQ approaches [28] have primarily focused on addressing the bimodal distribution of post-projection Key activations and the diverse distribution patterns of post-Softmax attention scores in SAM. Despite these efforts, existing methods fail to account for additional key factors crucial to quantization, resulting in severe performance degradation in ultra-low-bit scenarios. As shown in Table 1, existing methods exhibit significant precision loss at 5-bit quantization and collapse at 4-bit quantization. In this paper, we identify two additional challenges that profoundly affect SAM’s quantization performance, as illustrated in Fig. 1.
The first challenge arises from the heavy-tailed and skewed distribution of post-GELU activations. As depicted in Fig. 1(a), over 90% of activations are densely concentrated within -0.2 to 0, whereas a sparser but evenly distributed subset of large values, critical for inference accuracy, extends from 0 to 0.8. As illustrated in Fig. 2 (left panel), existing log2 and uniform quantizers fail to effectively handle this distribution. The log2 quantizer efficiently captures small, densely clustered values by allocating most of its quantization grids to this range. However, its exponentially spaced grid structure results in insufficient representation density for larger values, leading to high quantization errors in the upper range. On the other hand, the uniform quantizer provides consistent resolution across the full range, making it better suited for large, evenly spaced values, but it lacks sufficient resolution for small activations, introducing substantial quantization errors. This problem is further exacerbated as bit-width decreases, particularly in ultra-low-bit settings, where the limited number of grids imposes severe constraints. While non-uniform quantizers could offer a more suitable alternative, their hardware complexity makes deployment challenging [35, 3, 9]. To effectively address this issue, an efficient, hardware-compatible quantizer that captures both small and large activations is required.
The second challenge arises from high inter-channel variation, particularly in the activations of Query/Key/Value linear projections in the attention module and Linear projections in the MLP module [^1] A comprehensive analysis is provided in Appendix A.. These variations originate from LayerNorm operations and the unique activation distributions of the SAM mask decoder. As shown in Fig. 1(b), activation ranges in the Token-to-Image Value linear projection exhibit significant inter-channel disparities, making per-tensor quantization ineffective due to the challenge of finding a single optimal scale factor and zero point [1]. A large scale factor, selected to accommodate high-range channels, leads to reduced quantization granularity for low-range channels, often causing them to be rounded to zero. Conversely, a small scale factor, optimized for low-range channels, results in severe clipping in high-range channels, leading to significant information loss [43]. Moreover, while a zero point could compensate for activation skewness, the varying inter-channel median values prevent a single zero point from being optimal for all channels. While per-channel quantization can effectively mitigate quantization errors, it introduces considerable hardware inefficiencies. As depicted in Fig. 4, transferring per-channel quantization parameters between DRAM and compute units incurs a memory access overhead of up to several kB per layer [10, 36]. Storing these parameters on-chip can eliminate the transfer cost, but at the expense of a significantly larger memory footprint, requiring tens of thousands of on-chip registers, thereby increasing chip area utilization [^2] A detailed benchmarking is provided in Appendix B.. To address this trade-off, a granularity-aware quantization approach with hardware co-optimization is essential for balancing quantization accuracy and deployment efficiency.
3.3 AHCPTQ
To advance SAM quantization, we propose AHCPTQ, a framework specifically designed to overcome the two key challenges discussed earlier. As shown in Fig. 2, AHCPTQ leverages CAG and HLUQ, each targeting a specific quantization issue.
Figure 2: AHCPTQ framework: HLUQ refines quantization resolution for post-GELU activations, while CAG effectively groups parameters to manage channel-wise variations in linear projection activations to reduce quantization error.
3.3.1 Hybrid Log-Uniform Quantization
The heavy-tailed and skewed post-GELU activations hinder the quantization effectiveness of conventional uniform and log2 quantizers, while non-uniform quantizers, though potentially more adaptable, lack hardware efficiency. To enhance SAM quantization performance while ensuring hardware compatibility, we propose HLUQ, a method that reduces quantization errors in post-GELU activations by combining log2 and uniform quantization. Specifically, HLUQ applies log2 quantization to densely packed small values, where fine-grained precision is crucial, and uniform quantization to sparse, evenly distributed large values, ensuring minimal quantization error:
$x_{q}=\begin{cases}\text{clamp}\left(\left\lfloor-\log_{2}\frac{x}{s_{1}}\right\rceil,0,\hat{b}\right),&\text{if }x\leq s_{1},\\[5.0pt] \text{clamp}\left(\left\lfloor\frac{x-s_{1}}{s_{2}}\right\rceil,\hat{b},2^{k}-1\right),&\text{if }x>s_{1}.\end{cases}$ (6)
$x\approx\hat{x}=\begin{cases}s_{1}\cdot 2^{-x_{q}},&\text{if }x_{q}\leq\hat{b},\\ s_{2}\cdot x_{q}+s_{1},&\text{if }x_{q}>\hat{b}.\end{cases}$ (7)
Here, $s_{1}$ defines the log2 quantization scale, and $s_{2}$ determines the uniform quantization scale. The threshold $\hat{b}$ partitions the quantization grid, assigning log2 quantization to $[0,\hat{b}]$ and uniform quantization to $[\hat{b},2^{k}-1]$.
By adjusting $s_{1}$ and $\hat{b}$, HLUQ offers a high degree of adaptability, allowing it to accommodate various activation distributions. If $s_{1}$ spans the full range and $\hat{b}$ is close to $12^{k}-$ 1, HLUQ behaves like a log2 quantizer. In contrast, if $s_{1}$ and $\hat{b}$ are near zero, it essentially acts as a uniform quantizer, ensuring uniform step sizes. For heavy-tailed and skewed post-GELU activations, HLUQ can be configured with intermediate values of $s_{1}$ and $\hat{b}$, allowing it to capture small, densely distributed values using log2 quantization while retaining precision for large, sparse values through uniform quantization, as illustrated in Fig. 2 (left panel). Furthermore, HLUQ maintains the hardware efficiency of log2 and uniform quantization, introducing only minimal overhead to partition the input at $s_{1}$.
To initialize $\hat{b}$, $s_{1}$, and $s_{2}$, we introduce two auxiliary parameters, $\alpha$ and $\beta$, and search their optimal values during initial calibration by minimizing the following objective function:
$\mathop{\arg}\mathop{\min}_{\alpha,\beta}\mathbb{E}\left[\|\mathbf{X}\mathbf{W}-\hat{\mathbf{X}}\mathbf{W}\|_{F}^{2}\right]$ (8)
Here, $\alpha$ partitions the original activation range $r$, defining the quantization scales $s_{1}$ and $s_{2}$ as $s_{1}=\alpha\cdot r$ and $s_{2}=(1-\alpha)\cdot r$. Meanwhile, $\beta$ determines the allocation of quantization grids between log2 and uniform quantization segments, setting the threshold $\hat{b}$ as $\hat{b}=\beta\cdot(2^{k}-1)$. Once $\alpha$ and $\beta$ are obtained, the scales $s_{1}$, $s_{2}$, and the grid threshold $\hat{b}$ are configured to initialize HLUQ quantizer for subsequent reconstruction training.
3.3.2 Channel-Aware Grouping
The substantial inter-channel variation in activations presents a major obstacle for low-granularity per-tensor quantization, while fine-granularity per-channel quantization remains inefficient for hardware implementation. Interestingly, although linear projection activations exhibit substantial inter-channel variation, their statistical properties remain remarkably consistent across different samples. As shown in Fig. 3, the cosine similarity of normalized quantization parameters, searched over 100 samples per channel, is consistently close to 1.0, indicating that the optimal quantization parameters for each channel are largely invariant across samples. This stability presents an opportunity to leverage shared quantization parameters within grouped channels, improving quantization efficiency while maintaining accuracy.
Figure 3: Cosine similarity of normalized quantization parameter across 100 samples for each channel.
To balance quantization granularity with hardware efficiency, we propose CAG, which achieves accuracy comparable to per-channel quantization while significantly reducing parameter overhead, thereby enabling efficient on-chip deployment. The fundamental concept of CAG is to progressively group channels with similar activation distributions and assign them shared quantization parameters. As outlined in Algorithm 1, the process begins with quantization parameter initialization (scales and zero points) for each channel via model calibration, treating each channel as an independent group initially. Next, block-wise reconstruction is applied to refine both quantization parameters and weights, minimizing quantization error as formulated in Eq. 5, ensuring alignment with each group’s distribution characteristics. At designated milestones, channels with similar quantization parameters are progressively clustered, and the resulting centroids are adopted as shared quantization parameters for all channels within a group. This iterative process continues until the target group count is reached. The right panel of Fig. 2 illustrates an example of this grouping process for linear projection activations in Token-to-Image Value cross-attention, demonstrating how CAG effectively reduces the number of groups while preserving quantization performance. With a group number of 4, the edge accelerator can minimize quantization parameter overhead, reducing either data transmission or on-chip storage by 99.7%, as shown in Fig. 4. In AHCPTQ, we adopt on-chip storage for quantization parameters, requiring only 144 registers to store the scales and zero points of these 4 groups under 4-bit quantization. As depicted in Fig. 6, CAG maintains accuracy comparable to per-channel quantization, offering high hardware efficiency while significantly enhancing model performance.
Algorithm 1 Channel-Aware Grouping
0:  Total iterations $T$, milestones $t\in\{T_{1},T_{2},\dots,T_{J}\}$
0:  $K$ groups of optimized parameters ${\{s_{i},z_{i}\}}_{i=1}^{K}$
1:  Initialize quantization parameters ${\{s_{i},z_{i}\}}_{i=1}^{N}$ for $N$ channels, $G\leftarrow N$
2:  for $t=1,2,...,T$ do
3:     Update ${\{s_{i},z_{i}\}}_{i=1}^{G}$ according to Eq. 5
4:     if $t\in\{T_{1},T_{2},\dots,T_{J}\}$ then
5:        Apply K-Means to ${\{s_{i},z_{i}\}}_{i=1}^{G}$ to form centroids ${\{s_{i},z_{i}\}}_{i=1}^{G_{T_{j}}}$
6:        Assign each channel to the nearest centroid to form new groups
7:        Update number of groups $G\leftarrow G_{T_{j}}$
8:     end if
9:  end for
10:  return  ${\{s_{i},z_{i}\}}_{i=1}^{K}$
Figure 4: Hardware cost analysis of the linear projection layer in the SAM-H decoder under different quantization granularities.
3.4 Hardware Co-optimization
In this subsection, we briefly describe the strategy of hardware co-optimization for the proposed HLUQ and CAG. We provide a detailed implementation and illustration of the co-designed FPGA hardware in Appendix C.
HLUQ. To implement HLUQ, a dual processing-element (PE) architecture is employed to facilitate matrix multiplication within the HLUQ quantizer. The uniform quantization branch utilizes a standard integer multiplier and an adder tree, while the log2 quantization branch employs a bit-shifter and a decimal adder tree. The grid ratio between the two quantizers follows the $2^{-n}$ rule, ensuring that the MSBs of quantized values function as label bits. These label bits enable the accelerator to efficiently route quantized values to their respective PEs. Finally, the outputs from both branches are merged during dequantization, producing the final result.
CAG. To efficiently implement CAG, a small set of on-chip registers is used to store quantization parameters, and a customized quantizer logic is designed. Specifically, weights and activations are reordered by channel indices, ensuring that grouped channels are sequentially aligned. Weights are pre-processed offline, while activations are reordered on-chip. This reordering is integrated into the quantization and dequantization processes within linear projection computations. By employing counter-based logic for parameter switching, the design effectively minimizes hardware complexity while maintaining computational efficiency.
4 Experiments
4.1 Experiment and Implementation Details
We evaluate the effectiveness of AHCPTQ on instance segmentation using the mean Average Precision (mAP) metric on the COCO dataset[21]. The selected detectors include CNN-based Faster RCNN[31] and YOLOX[7], as well as Transformer-based H-Deformable-DETR[12] and DINO [40]. The bounding boxes generated by these detectors serve as prompts for SAM. For CNN-based detectors, the box threshold is set to 0.05, while Transformer-based detectors utilize a set of 100 adaptive anchors. Following the PTQ4SAM framework [28], we randomly sample 32 training images and use the MinMax approach to initialize quantization parameters. Block reconstruction is performed over 20,000 iterations on both attention and MLP blocks. To ensure fair comparisons [28, 23, 38, 39], we exclude the first and last layers or blocks from quantization while keeping all others quantized. HLUQ is applied to all Linear-2 activations in MLP blocks, with initial scales and grid thresholds determined by searching $\alpha\in\{0.1,0.3,0.5\}$ and $\beta\in\{\frac{1}{2},\frac{1}{4},\frac{1}{8}\}$. CAG is applied to all linear projection activations in Query/Key/Value projections of attention blocks and Linear-1 activations in MLP blocks, using a group number of 4. For other activations, we adopt per-tensor asymmetric quantization. For weights, we apply per-channel asymmetric quantization to maintain alignment with baseline settings.
4.2 Experimental Results
We evaluate AHCPTQ against baselines such as BRECQ [16], QDrop [38], and PTQ4SAM [28], with results summarized in Table 1 across four types of detectors. For SAM-L and SAM-H models, other baselines exhibit significant accuracy drops in the W5A5 configuration, while AHCPTQ reduces the mAP loss to within 2%. In the W4A4 configuration, where other baselines struggle to achieve basic functionality, AHCPTQ restores mAP to approximately 80% of the FP32 performance, particularly for SAM-H. For instance, AHCPTQ improves DINO’s mAP from 43.8% to 47.6% in the W5A5 configuration of SAM-H and from 2.3% to 36.6% in the W4A4 configuration of SAM-L, surpassing all baselines by a large margin. In the challenging SAM-B model, AHCPTQ doubles the mAP compared to PTQ4SAM in the W5A5 configuration, reduces the mAP loss to below 3% in W6A6, and restores mAP to a usable level in W4A4. For example, AHCPTQ improves H-DETR’s mAP by PTQ4SAM from 2.8% to 14.1%, 16.9% to 32.1%, and 30.7% to 36.6% in W4A4, W5A5, and W6A6 configurations, respectively. These results indicate that AHCPTQ consistently surpasses baselines, addressing the challenges identified in Sec. 3.2 and advancing SAM quantization to lower bit-widths with superior effectiveness.
Table 1: Quantization results of AHCPTQ for instance segmentation on the COCO dataset across four detector types. Floating-point (FP32) mAP values for SAM-B/L/H are shown below each detector name for reference. AHCPTQ achieves a new state-of-the-art performance, outperforming all existing baselines.
Detector Method SAM-B SAM-L SAM-H
FP32: (B/L/H) W4A4 W5A5 W6A6 W4A4 W5A5 W6A6 W4A4 W5A5 W6A6
Faster R-CNN BRECQ 0.2 16.7 28.0 5.0 31.8 35.2 17.5 31.3 35.8
QDrop 2.3 12.9 26.2 0.8 31.9 35.0 6.0 32.6 36.0
33.1/36.0/36.8 PTQ4SAM 2.7 14.4 26.8 2.4 33.0 35.5 6.7 33.3 36.2
AHCPTQ 11.7 27.5 31.6 27.4 34.8 35.6 31.4 35.7 36.3
YOLOX BRECQ 0.2 19.0 31.9 6.3 35.3 39.4 19.7 34.7 39.7
QDrop 2.6 15.6 30.3 1.0 36.2 39.4 6.8 36.0 40.1
37.2/40.4/41.0 PTQ4SAM 3.8 18.4 30.9 2.4 37.1 39.9 7.4 37.1 40.3
AHCPTQ 13.4 31.8 35.4 31.0 39.1 40.0 35.2 40.0 40.4
H-DETR BRECQ 0.3 11.2 32.0 5.2 36.1 40.4 19.1 35.3 40.6
QDrop 2.0 13.1 30.5 1.3 37.0 40.3 6.9 37.0 41.1
38.2/41.5/42.0 PTQ4SAM 2.8 16.9 30.7 2.6 38.1 40.9 7.1 38.0 41.4
AHCPTQ 14.1 32.1 36.6 32.3 40.3 41.0 35.6 40.9 41.5
DINO BRECQ 0.2 13.5 34.8 3.6 41.4 47.0 20.7 40.3 47.2
QDrop 1.9 13.4 34.5 1.0 42.7 47.0 7.0 42.4 47.9
44.5/48.6/49.1 PTQ4SAM 1.9 17.6 35.1 2.3 44.1 47.8 8.9 43.8 48.2
AHCPTQ 16.8 36.7 41.9 36.6 46.8 47.9 41.2 47.6 48.3
Note: This paper reports the corrected performance of PTQ4SAM and QDrop after fixing a bug (refer to code repo.) in the PTQ4SAM [28] framework.
4.3 Ablation Studies
Ablation of Components. Table 2 details the ablation results for AHCPTQ, which are tested on YOLOX and H-Deformable-DETR in the W4A4 configuration. The results demonstrate that each technique plays a vital role in enhancing SAM’s performance, with the highest gains achieved when both CAG and HLUQ are applied together. For SAM-B and SAM-L models, inter-channel variance emerges as the primary challenge. Thus, while applying HLUQ alone leads to moderate improvements, applying CAG yields significant performance gains. Once the major challenge is addressed by CAG, HLUQ becomes essential for further improvement. For instance, HLUQ achieves an additional 8.3% mAP improvement in the SAM-L model of YOLOX when CAG is applied. For SAM-H, heavy-tailed and skewed activations become the dominant limitation. As a result, HLUQ contributes more significantly than CAG, and their combined application results in a 28% mAP improvement.
Ablation of Quantizers. To investigate the effectiveness of HLUQ, we analyze SAM performance by substituting HLUQ with uniform and log2 quantizers, respectively. As the standard log2 quantizer cannot handle the negative branch of post-GELU activations, a floating-point bias is added for reasonable comparison. The results for YOLOX and H-Deformable-DETR under the W4A4 configuration are presented in Fig. 5. HLUQ significantly outperforms the uniform quantizer and surpasses the biased log2 quantizer, demonstrating its ability to improve resolution under low-bit quantization and reduce quantization error, thereby boosting SAM performance.
(a) YOLOX
(b) H-DETR
Figure 5: Ablation study comparing the effectiveness of HLUQ with other quantizers in the W4A4 configuration.
Dependence on Group Number. To analyze the effect of group number on SAM’s performance in CAG, we evaluate mAP as the group number varies from 2 to 32 in YOLOX and H-Deformable-DETR under the W4A4 configuration. The results, presented in Fig. 6, also include per-tensor and per-channel configurations for reference. For SAM-L and SAM-H models, mAP increases sharply and saturates at a group number of 2, with slight fluctuations and marginal improvements closer to the per-channel configuration. For the SAM-B model, the turning point starts at a group number of 4, with a larger performance gap of approximately 3% compared to the per-channel configuration. Considering the need for a consistent configuration across models while minimizing hardware overhead, we choose a group number of 4 in the AHCPT framework.
Table 2: Results of the ablation study analyzing the contribution of components in the W4A4 configuration.
Detector CAG HLUQ S.-B S.-L S.-H
YOLOX $\times$ $\times$ 3.7 2.5 7.1
$\checkmark$ $\times$ 10.8 22.7 20.2
$\times$ $\checkmark$ 5.7 5.7 24.8
$\checkmark$ $\checkmark$ 13.4 31.0 35.2
H-DETR $\times$ $\times$ 3.1 2.7 7.3
$\checkmark$ $\times$ 11.2 22.1 20.5
$\times$ $\checkmark$ 4.1 6.5 24.6
$\checkmark$ $\checkmark$ 14.1 32.3 35.6
(a) YOLOX
(b) H-DETR
Figure 6: Dependence of SAM performance on group number in CAG for the W4A4 configuration.
4.4 Hardware Validation
To evaluate the resource efficiency and practical performance of AHCPTQ in real-world applications, we developed an FPGA accelerator for executing the MLP block in the SAM mask decoder. CAG is applied to Linear-1 activations to mitigate channel-wise variation, while HLUQ is applied to Linear-2 activations to accommodate heavy-tailed and skewed post-GELU distributions. The accelerator is designed in Verilog, synthesized using Vivado Design Suite, and implemented on a Xilinx Artix UltraScale+ FPGA running at 200 MHz. The overall system consists of an FPGA accelerator for computation, a DDR4 DRAM for data buffering, and a host PC for transferring activations via Ethernet. In accordance with Sec. 3.4, the accelerator uses MSBs indexing to assign activations to appropriate PEs—either multipliers or bit-shifters—for HLUQ. We fuse the activation reordering logic into the quantization/dequantization steps for CAG and pipeline the quantization process with activation functions, effectively eliminating extra quantization overhead in HLUQ. To establish a comparative benchmark, we implemented two baselines: a standard FP32 accelerator and a default 8-bit integer (INT8) accelerator. Floating-point operations, including multiply-and-accumulate in the FP32 accelerator and quantization/dequantization process in the integer accelerator, were implemented using the Floating-Point Operation IP generator with on-chip DSP resources. As shown in Table 3, the FP32 accelerator is limited by the number of available on-chip DSP resources, supporting only 8 parallel PE lanes. In contrast, both the default INT8 and AHCPTQ INT4 accelerators replace DSP-based PEs with LUT-based PEs, allowing for a significant increase in parallelism to 32 and 64 lanes, respectively. Meanwhile, DSP resources in integer accelerators are allocated to quantization and dequantization processing, optimizing overall resource utilization for improved system-level performance. As a result, 4-bit AHCPTQ reduces the computational complexity and overhead of data movement considerably, thereby achieving 8.64×8.64$\times$8.64 × gains in energy efficiency and 7.89×7.89$\times$7.89 × speedup over its floating-point counterpart. This demonstrates strong potential for efficient SAM deployment on edge devices.
Table 3: Analysis of resource utilization and system-level performance of AHCPTQ on an FPGA platform.
FP32 INT8 AHCPTQ INT4
LUT Usage 19,071 43,015 34,199
DSP Usage 447 153 213
Parallelism 8 32 64
GOPS/W 5.04 21.97 43.57
Speedup 1.00×1.$0$ 3.96×3.96$\times$3.96 × 7.89×7.89$\times$7.89 ×
5 Limitation
This study focuses on evaluating the feasibility of CAG and HLUQ on FPGA and ASIC platforms. We expect that similar enhancements in energy efficiency and speedup could be realized on CPU and GPU platforms due to the reduction in FLOPs and data movement overhead. The deployment of AHCPTQ on these platforms is left for future research.
6 Conclusion
This paper presents AHCPTQ, a novel PTQ framework for SAM, designed to efficiently handle complex activation distributions while ensuring compatibility with hardware acceleration. AHCPTQ incorporates HLUQ to manage heavy-tailed and skewed activations, and CAG to address high inter-channel variation, significantly improving quantization performance. Extensive experiments on multiple SAM variants and quantization settings demonstrate its superior performance, establishing a new state-of-the-art in SAM quantization. Moreover, FPGA-based evaluations confirm its real-world deployability with substantial speedup and energy efficiency gain, providing valuable insights for practical applications.
Supplementary Material
Appendix A Analysis of Inter-Channel Variation and Inter-Sample Similarity in SAM Model
In this section, we provide an in-depth analysis of inter-channel variation and inter-sample similarity using the SAM-B model with YOLOX as the prompt detector. We extract the quantization ranges for channel indices 50 to 100 to examine channel-wise variation across multiple layers where CAG is applied. Additionally, we determine the optimal scale and zero point for each channel using different samples to assess parameter consistency across 100 samples.
In the image encoder, Figs. 7(a) and 8(a) show that QKV projection and Linear-1 activations in the MLP block exhibit channel-wise variation due to LayerNorm operations.
(a) Range distribution for channel indices 50 to 100.
(b) Cosine similarity across 100 samples for each channel.
Figure 7: Range distribution and cosine similarity of QKV projection activations in the image encoder.
(a) Range distribution for channel indices 50 to 100.
(b) Cosine similarity across 100 samples for each channel.
Figure 8: Range distribution and cosine similarity of Linear-1 activations in the MLP block of the image encoder.
In the mask decoder, channel-wise variation in certain layers poses a major challenge for low-bit quantization. In Token-to-Image cross-attention, linear projection activations in the Query and Value projections serve as the primary challenges, as shown in Fig. Figs. 9(a) and 1(b). Notably, activation outliers in the Query projection extend the quantization range to approximately 400, severely increasing quantization error under per-tensor quantization.
(a) Range distribution for channel indices 50 to 100.
(b) Cosine similarity across 100 samples for each channel.
Figure 9: Range distribution and cosine similarity of pre-projection activations for Token-to-Image Query in the mask decoder.
In Image-to-Token cross-attention, linear projection activations in the Key and Value projections exhibit distribution variations that degrade SAM’s performance. As shown in Figs. 10(a) and 11(a), Key activations remain relatively stable, whereas Value activations demonstrate higher variance.
(a) Range distribution for channel indices 50 to 100.
(b) Cosine similarity across 100 samples for each channel.
Figure 10: Range distribution and cosine similarity of pre-projection activations for Image-to-Token Key in the mask decoder.
(a) Range distribution for channel indices 50 to 100.
(b) Cosine similarity across 100 samples for each channel.
Figure 11: Range distribution and cosine similarity of pre-projection activations for Image-to-Token Value in the mask decoder.
Lastly, the range distribution of Linear-1 activations in the MLP block of the mask decoder is shown in Fig. 12(a).
(a) Range distribution for channel indices 50 to 100.
(b) Cosine similarity across 100 samples for each channel.
Figure 12: Range distribution and cosine similarity of Linear-1 activations in the MLP block of the mask decoder.
Fortunately, we verified that in all layers exhibiting severe channel-wise variation, the optimal quantization parameters remain stable across different samples. As shown in Figs. 7(b), 8(b), 9(b), 3, 10(b), 11(b) and 12(b), the normalized cosine similarity scores remain consistently high across most channels. Therefore, the CAG-based grouping strategy proves to be a highly reliable approach for enhancing SAM’s quantization performance.
Appendix B Hardware Cost Benchmark Across Different Quantization Granularity
To benchmark the hardware cost associated with different quantization granularities, we analyze the implementation overhead of the SAM-H model under three configurations: per-tensor quantization, per-channel quantization, and CAG in AHCPTQ. Deep learning accelerators typically handle quantization parameters in two ways:
(1) Storing scales and zero points in on-chip registers. This approach completely eliminates data transfer between the off-chip DRAM and the accelerator, ensuring zero transmission latency and cost. However, it increases area and resource overhead due to the large number of registers required.
(2) Storing scales and zero points in off-chip DRAM. This method reduces the on-chip resource overhead but incurs substantial energy and latency costs due to frequent DRAM accesses [10].
As illustrated in Fig. 4, any combination between these two approaches follows a trade-off curve—reducing one overhead inevitably increases the other. To mitigate this trade-off, we propose CAG, which clusters quantization parameters to significantly reduce their count. By grouping scale and zero-point values into just four clusters, CAG reduces the required registers in (1) by 99.7% or the data transfer cost in (2) by 99.7%. This approach enables hardware efficiency close to per-tensor quantization while maintaining accuracy comparable to per-channel quantization. Since only 144 registers are needed to store four groups of scale and zero points in 4-bit quantization, we opt to store them on-chip, ensuring a more efficient and simplified hardware design.
Appendix C Hardware Implementation Details
Our AHCPTQ approach is implemented on an FPGA, with a custom accelerator designed to validate the quantization strategy. Fig. 13(a) depicts the validation system, which includes a host PC for data transfer via Ethernet, DDR4 DRAM for data buffering, and a Xilinx Artix UltraScale+ FPGA configured with the accelerator. The general architecture of the accelerator, shown in Fig. 13(b), incorporates four groups of buffers that leverage on-chip BRAM resources. These buffers temporarily store input activations, model weights, and output activations before and after quantization. The accelerator supports two configurations of PEs: (1) an 8-input integer multiplier-and-adder PE for uniform quantization, where both inputs and outputs are integers and partial sums are updated using an accumulator, and (2) an 8-input decimal bit-shift-and-adder PE for log2 quantization, with integer inputs, decimal outputs, and a decimal accumulator for partial sums. In our 4-bit AHCPTQ implementation, the accelerator uses 64 lanes for multiplier and bit-shift PEs. To enable HLUQ, weights and activations must be dynamically allocated to their corresponding PEs. This allocation logic is embedded in the controller, which indexes the MSBs of activations. When data is read from the IA buffer, weights and activations are automatically distributed to the corresponding PE types.
(a) Evaluation system.
(b) Accelerator architecture.
Figure 13: Overview of the evaluation system and the accelerator architecture for the AHCPTQ configuration.
After completing the integer and decimal inner product operations, the output values are transferred to the quantization processor in the FP32 domain. A small set of quantization parameter registers loads the scales required for the current and next layers, switching addresses to provide these scales to the quantization processor. The output activations are then fed into the dequantization unit, where, if HLUQ is applied, the uniform and log2 branches are merged. Subsequently, the activations pass through the activation functions and quantization units, ensuring the resulting integer values can be directly used by the next layer. Finally, the activations are sent to the output buffer. The dequantization, activation function, and quantization processes are designed in a pipeline to minimize latency, which is critical for HLUQ deployment in practical applications.
Weights are grouped offline and reordered based on group indices, while activations are reordered on-the-fly. At the start of layer computation, the accelerator accesses the DRAM to transfer weights and activations to the weight and IA buffers. Simultaneously, the quantization processor loads the quantization parameters required for quantization and dequantization. The controller then dispatches paired weights and activations to the appropriate PE lanes based on the MSB of the activations. For instance, with $\beta=\frac{1}{2}$ as described in Sec. 3.3.1, activations with an MSB of ’1’ are routed to multiplier PE lanes, while those with an MSB of ’0’ are sent to bit-shift PE lanes. Once computation is completed, the results stored in the accumulators are clustered into two categories and passed to their respective uniform or log2 dequantization units. The dequantized FP32 activations are then written back to the quantization buffer using a tailored addressing logic to restore their default sequence. Following this, the activations undergo the activation function and quantization for the next layer before being sent to the output buffer and ultimately recycled back to DRAM. To streamline the process, the reorder logic is fused into the DRAM controller, which writes activations to grouped addresses in DRAM when CAG is applied to the next layer.
Figure 14: FPGA validation environment.
We implemented the RTL of the accelerator, Ethernet interface, and DRAM controller in Verilog, synthesized the design using Vivado Design Suite, and deployed it on the ALINX AXAU15 development board. As illustrated in Fig. 14, weights and activations are stored in on-board DDR4 DRAM and transferred via Ethernet from the host PC. The accelerator operates at 200 MHz to assess speedup and energy efficiency. For comparison, we implemented two baseline accelerators alongside our AHCPTQ configuration: (1) a standard FP32 implementation, (2) a default INT8 implementation. In all designs, floating-point operations such as quantization and dequantization are handled by an IP generator utilizing on-chip DSP resources. The detailed experimental results are presented in Sec. 4.4.
Appendix D Experiment on Vision Transformers
To ensure that AHCPTQ generalizes to other vision models and tasks, we evaluate its effectiveness on ImageNet for image classification using DeiT. We integrate AHCPTQ into I&S-ViT [44], the latest state-of-the-art PTQ framework, by replacing its original post-GELU quantizer with HLUQ.
Table 4: Comparison of W4A4 PTQ methods on DeiT based on image classification accuracy on the ImageNet dataset.
Method Opti. DeiT-T DeiT-S DeiT-B
FQ-ViT [22] $\times$ 0.10 0.10 0.10
PTQ4ViT [39] $\times$ 36.96 34.08 64.39
APQ-ViT [4] $\times$ 47.94 43.55 67.48
BRECQ [16] $\checkmark$ 55.63 63.73 72.31
QDrop [38] $\checkmark$ 61.93 68.27 72.60
PD-Quant [23] $\checkmark$ 62.46 71.21 73.76
RepQ-ViT [20] $\times$ 57.43 69.03 75.61
I&S-ViT [44] $\checkmark$ 65.21 75.81 79.97
AHCPTQ $\checkmark$ 66.11 76.12 80.07
In vision transformers (ViTs), inter-channel variation can be effectively addressed by reparameterizing LayerNorm’s weight and bias, as LayerNorm consistently precedes the QKV linear projection. Consequently, we follow the settings of RepQ-ViT [20] and I&S-ViT [44], applying reparameterization to reduce inter-channel variance, making CAG unnecessary in this case. However, in SAM’s decoder, the LayerNorm placement differs significantly from ViT’s image encoder, making efficient reparameterization infeasible. This fundamental difference motivated the introduction of CAG as a dedicated PTQ solution for SAM with high hardware efficiency.
We perform PTQ on DeiT-T, DeiT-S, and DeiT-B, comparing AHCPTQ against static PTQ methods (FQ-ViT [22], PTQ4ViT [39], APQ-ViT [4], RepQ-ViT [20]) and optimization-based PTQ methods (BRECQ [16], QDrop [38], PD-Quant [23], I&S-ViT [44]). As shown in Table 4, AHCPTQ achieves the highest classification accuracy in W4A4 quantization, surpassing all competing methods.